OpenStudy (anonymous):

1 down, 3 more to go. Make "h" the subject of A = pir^2 +2pirh (yes, that is pi * "r squared" PLUS 2pi * r * h)

OpenStudy (anonymous):

lol u are never going to learn if u get others to do all the questions :|

OpenStudy (anonymous):

\[A=\pi r^2 +2 \pi r h\] solve for h \[A-\pi r^2=2\pi rh\] \[\frac{A-\pi r^2}{2\pi r}=h\]

OpenStudy (anonymous):

area of cylinder

OpenStudy (anonymous):

@elecengineer "cheat in learning"

OpenStudy (anonymous):

At this point, I'm on a timed limit here. I have no time to debate whether or not I want to actually learn OR get it handed down. I need this in tomorrow.

OpenStudy (anonymous):

lol